OpenStudy (anonymous):

Divide. 3\5 \ 1\4 A. 3\20 b.5\12 c.4\9 d.2 2\5

OpenStudy (anonymous):

@LegendarySadist @Hero

OpenStudy (anonymous):

When you divide fractions, you multiply the first fraction by the reciprocal of the second.

OpenStudy (anonymous):

Is it A?

OpenStudy (anonymous):

No, it would not be A. Remember \[\large \sf \frac{a}{b} \div \frac{c}{d}~=~\frac{a}{b} \times \frac{d}{c}\]

OpenStudy (anonymous):

d?

OpenStudy (anonymous):

Correct :)